# Towards Adapting Open-Source Large Language Models for Expert-Level Clinical Note Generation

### v5

@@ Abstract
Proprietary Large Language Models (LLMs) such as GPT-4 and Gemini have demonstrated promising capabilities in clinical text summarization tasks. However, due to patient data privacy concerns and computational costs, many healthcare providers prefer using small, locally-hosted models over external generic LLMs. This study presents a comprehensive domain- and task-specific adaptation process for the open-source LLaMA-2 13 billion parameter model, enabling it to generate high-quality clinical notes from outpatient patient-doctor dialogues. Our process incorporates continued pre-training, supervised fine-tuning, and reinforcement learning from both AI and human feedback. We introduced a new approach, DistillDirect, for performing on-policy reinforcement learning with Gemini 1.0 Pro as the teacher model. Our resulting model, LLaMA-Clinic, can generate clinical notes comparable in quality to those authored by physicians. In a blinded physician reader study, the majority (90.4%) of individual evaluations rated the notes generated by LLaMA-Clinic as “acceptable” or higher across all three criteria: real-world readiness, completeness, and accuracy. In the more challenging “Assessment and Plan” section, LLaMA-Clinic scored higher (4.2/5) in real-world readiness than physician-authored notes (4.1/5). We highlight key considerations for future clinical note-generation tasks, emphasizing the importance of pre-defining a best-practice note format, rather than relying on LLMs to determine this for clinical practice.
@@
Anonymous ACL submission
@@ 1 Introduction
Recent advancements in LLMs have transformed the field of natural language processing (NLP). However, the application of LLMs in the medical domain is still in its early stages  he2023survey; zhou2023survey. Proprietary LLMs, such as GPT-4 and Med-PaLM, have demonstrated impressive capabilities in medical knowledge and clinical NLP tasks nori2023capabilities; singhal2023expertlevel; van2023clinical. However, most proprietary LLMs have limited flexibility for domain-specific fine-tuning, primarily due to restricted access to model weights. Additionally, proprietary LLMs raise several concerns pertinent to the healthcare sector, including HIPAA compliance, data security, cost, and transparency of training data marks2023ai; clusmann2023future; Adimi2023; nyt_article_2023.
@@
The emergence of powerful open-source LLMs has opened up opportunities for domain-specific fine-tuning within the clinical field, yielding promising results han2023medalpaca; wang2023huatuo; li2023chatdoctor; wu2023pmcllama; toma2023clinical; chen2023meditron. However, most research on open-source models has concentrated on medical knowledge injection rather than practical applications in real-world clinical workflow such as clinical note generation.
@@
In this work, we address a practical question clinicians face in their everyday routine: How can we best adapt an open-source LLM for the specific use case of clinical note generation? Clinical note documentation represents a significant burden for healthcare practitioners ammenwerth2009time and appears to be a natural application for LLMs, given their remarkable generative capabilities lee2023benefits. Recent research on LLMs for clinical text summarization found that LLM-generated outputs are preferred over human summaries for their completeness and accuracy van2023clinical.
@@
Figure 1: Overview of Study Design. We conducted a comprehensive domain- and task-specific adaptation process for the LLaMA-2-13B model. This process included continued pre-training, supervised fine-tuning, and reinforcement learning from AI and human feedback. Finally, we evaluated our model’s outputs against those created by physicians and Gemini Pro through a blinded expert evaluation. We used Gemini 1.0 Pro as the teacher model in this study.
@@
The 2023 ACL ClinicalNLP and CLEFImage workshops explored the generation of clinical notes from patient-doctor conversations using the newly released Ambient Clinical Intelligence Benchmark (ACI-BENCH) dataset abacha2023overview; yim2023overview. The most notable results were achieved using GPT-4 along with few-shot in-context learning giorgi2023wanglab; van2023clinical. However, these initial exploratory studies for clinical note generation with LLMs revealed significant limitations:
@@
  • Issue with metrics: The evaluation relied solely on automatic metrics of lexical similarity to the “reference” notes, such as ROUGE scores, which may not accurately reflect human preferences moramarco2022human; van2023clinical.
@@
  • Limited fine-tuning datasets: Previous studies conducted only limited supervised fine-tuning (SFT) of open-source LLMs using small training datasets, thereby not fully exploring their potential for domain- and task-specific adaptation abacha2023overview; yim2023overview; van2023clinical.
@@
    Clinical LLM fine-tuning playbook: We provide a comprehensive guide for healthcare organizations to fine-tune local LLMs on their own data. Our relatively small, open-source LLM achieved expert-level performance, scoring higher (4.2/5) in real-world readiness compared to physician-authored notes (4.1/5) in the “Assessment and Plan” section, as evaluated by clinical experts.
@@
    Key considerations for future clinical note-generation tasks: We underscored the importance of pre-defining a best-practice note format, rather than relying on LLMs to determine this for clinical practice. We also recommend a workflow for physicians to work with AI-generated notes in real-world practice, ensuring physicians’ full oversight and ultimate accountability for the AI-generated content.
@@ 2.1 Problem Formation
Given a recorded dialogue from a patient-doctor clinic encounter, we task LLMs to generate a high-quality outpatient note akin to one written by a clinician. This scenario is becoming increasingly prevalent due to the rising popularity of both in-person and virtual scribes, and automatic speech recognition devices in real-world settings yim2023aci. Our research focus is on generating the “Subjective” and “Assessment and Plan” sections of outpatient notes. This decision is based on feedback from our physician coauthors, which indicates that discussing all details of the “Objective” section, such as physical examination results, is impractical during actual clinical encounters. Furthermore, much of the objective data, including lab and imaging results, are directly integrated into Electronic Medical Records (EMR), making clinical notes generation for those sections easy and sometimes unnecessary.
@@ 2.2 ACI-BENCH
ACI-BENCH represents the largest clinic dialogue-note dataset publicly available to date, comprising 207 cases yim2023aci. The dataset’s dialogues were crafted by a team with medical expertise, and its clinical notes were initially generated using an automatic note generation system, then reviewed and revised by domain experts yim2023aci. The ACI-BENCH dataset was previously utilized to benchmark the performance of outpatient note generation systems using automatic metrics that evaluate lexical similarity abacha2023overview; yim2023overview. Importantly, we observed notable variation in the format, style, and quality of the “reference notes” within ACI-BENCH, especially in the section of “Assessment and Plan”. While this diversity mirrors the reality of clinical practice, where different doctors may produce vastly different notes, it presents a challenge to use these notes as a “gold standard” for training an LLM to replicate.
@@ 2.3 Distilled DPO
Given the significant time and financial costs associated with collecting preference data, utilizing pre-collected preference datasets, such as those employed in distilled DPO, is a common practice. Distilled DPO involves generating a collection of responses for each prompt from various LLMs (Figure 2A) tunstall2023zephyr. These responses are then evaluated by a teacher model (e.g., GPT-4) to provide preference feedback. Applying reinforcement learning with AI feedback (RLAIF) with distilled DPO has yielded encouraging outcomes tunstall2023zephyr. However, distilled DPO’s reliance on a pre-collected preference dataset renders it suboptimal due to the off-policy and offline characteristics.
@@ 3 DistillDirect
3.1 Online vs. Offline Training and On-policy vs. Off-policy Training
@@ 3.2 RLAIF with DistillDirect
In our study, we employed RLAIF with DistillDirect, followed by Reinforcement Learning from Human Feedback (RLHF), which provides several advantages. First, for each iteration of DistillDirect, we used the target policy’s outputs as “reject samples” to inform feedback, ensuring on-policy learning as opposed to the off-policy training in previous work tunstall2023zephyr. Second, during the RLHF phase, we gathered human preferences on responses generated by the target policy, promoting online and on-policy training (Figure 2C).
@@ 4.1 Experimental Design
Model Selection: We selected Meta’s LLaMA-2 model with 13 billion parameters (LLaMA-2-13B) and conducted experiments using both the base and chat models touvron2023llama. Gemini 1.0 Pro (hereafter referred to as Gemini Pro) by Google was selected as the teacher model for generating reference notes team2023gemini.
Experiments Overview: We demonstrated the experiments pipeline in Figure 1. First, we undertook domain-specific adaptation of LLaMA-2-13B through continued pretraining on discharge summaries from MIMIC-IV. This was followed by task-specific fine-tuning with SFT through instruction tuning. Next, we conducted RLAIF using DistillDirect, our enhanced approach to performing DPO on a distilled dataset. We then selected the model that performed best after SFT and RLAIF, as optimized by the ROUGE-1 score against reference notes, for RLHF through DPO. Finally, a panel of physicians conducted a blinded evaluation of the notes authored by our LLaMA-Clinic model, Gemini Pro, and other physicians. All training was performed using low-rank adaptation (LoRA).
@@ 4.2 Dataset and Preprocessing
Modified ACI-BENCH: Given above mentioned limitation of ACI-BENCH, we established a simple yet specific note format, recognized as “best practice” by a panel of licensed internal medicine physicians, to standardize our training approach. For model training, we retained only the dialogue section from ACI-BENCH and employed Gemini Pro to generate notes based on the “best practice” format, serving as our reference notes. We demonstrated two examples of clinical notes before and after the change in Appendix Figure 11.
Dialogue-G: We created a synthetic dataset of clinical dialogue-note pairs using Gemini Pro. This dataset, which we have named Dialogue-G, comprises 1,291 cases. We compiled transcribed outpatient notes from the publicly available synthetic MTSamples dataset MTSamples; hu2024improving and utilized Gemini Pro to transform these notes into dialogues. Subsequently, we used these dialogues as inputs for Gemini Pro once again, this time to generate clinical notes based on the “best practice” format described above.
MIMIC-IV: MIMIC-IV encompasses 431,231 unique hospital admissions from Beth Israel Deaconess Medical Center in Boston, Massachusetts johnson2023mimic. We utilized discharge summaries from MIMIC-IV for continued pretraining. Notably, the “brief hospital course” section of the discharge summaries is structurally akin to the “assessment and plan” section in outpatient notes. We compiled a subset of discharge summaries with only the “brief hospital course” using methods detailed in wang2024drg, referred to as “Discharge-short”. We denoted the complete discharge summaries dataset as “Discharge-long”.
@@ 4.3 Experiment setup
Dataset Split for SFT, RLAIF and RLHF: We combined the training subsets from ACI-BENCH (dialogue n = 67) and Dialogue-G (dialogue n = 1291), then split this data equally for SFT and RLAIF, stratified by the data source. For RLHF, we utilized dialogues from the training, task2, and task3 subsets of ACI-BENCH (dialogue n = 147) as input prompts and sampled from our model’s outputs for human preference labeling.
@@
Physician Reader Study: The three internal medicine physicians engaged in preference data collection were tasked with writing clinical notes based on conversations from the ACI-BENCH test1 subset, adhering to the pre-defined “best practice” format. The physician-authored notes, alongside those generated by Gemini Pro and LLaMA-Clinic, were reviewed by three additional physicians, who were not involved in the preference labeling.
@@ 5.1 Analysis of Continued Pretraining
We presented the training loss curve in Figure 3. Across all experiments, the training loss rapidly decreased after the initial few hundred steps, then leveled off, showing minimal improvement thereafter. The lowest training loss achieved with the Discharge-long dataset is approximately 0.9, whereas, with the Discharge-short dataset, it remained around 1.4. The trajectories of training loss were similar for both the chat and base models. When experimenting with various training strategies and hyperparameters, we frequently observed training loss spikes that were slow to recover, as shown in Appendix Figure 5. We proceeded with models that did not exhibit loss spikes for further studies (i.e., our SFT and RLAIF experiments).
@@ 5.2 Analysis of SFT and RLAIF
Models CP SFT DistillDirect CP SFT DistillDirect
Models underwent continued pretraining, SFT and RLAIF
 13B_short 0.1488 0.2463 0.4775 0.0899 0.2277 0.4956
 13B_long 0.1612 0.2525 0.4494 0.1289 0.2565 0.4578
 13B_chat_short 0.1520 0.3475 0.4878 0.1162 0.3055 0.5182
 13B_chat_long 0.1708 0.3463 0.4601 0.1536 0.3620 0.4662
Models underwent SFT and RLAIF without continued pretraining
 13B - 0.2813 0.4994 - 0.2977 0.4941
 13B_chat - 0.2874 0.4783 - 0.3301 0.4893
Out-of-box models*
 13B 0.0329 0.0100
 13B_chat 0.3585 0.4543
@@
Table 1: ROUGE-1 Scores after SFT and RLAIF. All tests performed on validation subset of modified ACI-BENCH. Models including “chat” represent LLaMA-2 chat models, and otherwise represent LLaMA-2 base models. Models ended in “long” were pretrained using Discharge-long dataset, while models ended in “short” were pretrained using Discharge-short dataset. Bolded scores denote the best performance. A higher ROUGE-1 score indicates better performance. The out-of-box chat model outperformed the base model. Continued pretraining with MIMIC-IV discharge summaries initially led to performance deterioration, while SFT and DistillDirect significantly enhanced performance of all models. *: Out-of-box models did not undergo CP, SFT or RLAIF. CP: continued pretraining. RLAIF: reinforcement learning from AI feedback. SFT: supervised fine-tuning.
@@
As the primary objective of SFT and RLAIF is to align the output of LLaMA-2 with that of Gemini Pro, we evaluated the ROUGE-1 score, a measure of lexical similarity, against reference notes created by Gemini Pro. A higher ROUGE-1 score indicates better performance, with scores above 0.5 generally considered good abacha2023overview; yim2023overview. We reported the ROUGE-1 scores post-SFT and RLAIF, alongside those from the unmodified LLaMA-2, in Table 1. As anticipated, the vanilla chat model outperformed the base model out-of-box. However, continued pretraining with out-of-distribution corpora (i.e., MIMIC-IV discharge summaries) compromised the chat model’s capacity to follow instructions. SFT significantly enhanced performance, particularly for the chat model over the base model. The model’s performance was further boosted by RLAIF using DistillDirect. Notably, our training with DistillDirect frequently encountered instability—a well-known challenge in reinforcement learning ding2020challenges. With carefully selected training hyperparameters, DistillDirect could achieve performance improvements without overfitting, despite the limited training data and repetitive use of the same prompts in each training cycle (see Appendix Table 5). We have detailed our experiments, including ablation studies, to find a stable training setup in Appendix Method A.3.
@@
When comparing models pretrained with Discharge-long versus Discharge-short, the latter consistently scored higher despite a higher training loss in the pretraining phase. The chat model pretrained with Discharge-short emerged as the top performer, and we selected this model checkpoint for the final RLHF stage. Interestingly, the vanilla LLaMA-13B models, directly subjected to SFT and RLAIF without prior continued pretraining, exhibited strong performances, including the highest score for “Subjective”.

### v6

@@ Abstract
Proprietary Large Language Models (LLMs) such as GPT-4 and Gemini have demonstrated promising capabilities in clinical text summarization tasks. However, due to patient data privacy concerns and computational costs, many healthcare providers prefer using small, locally-hosted models over external generic LLMs. This study presents a comprehensive domain- and task-specific adaptation process for the open-source LLaMA-2 13 billion parameter model, enabling it to generate high-quality clinical notes from outpatient patient-doctor dialogues. Our process incorporates continued pretraining, supervised fine-tuning, and reinforcement learning from both AI and human feedback. We introduced a new approach, DistillDirect, for performing on-policy reinforcement learning with Gemini 1.0 Pro as the teacher model. Our resulting model, LLaMA-Clinic, can generate clinical notes comparable in quality to those authored by physicians. In a blinded physician reader study, the majority (92.8%) of individual evaluations rated the notes generated by LLaMA-Clinic as “acceptable” or higher across three criteria: real-world readiness, completeness, and accuracy. In the more challenging “Assessment and Plan” section, LLaMA-Clinic matched physician-authored notes in real-world readiness score. We highlight key considerations for future clinical note-generation tasks, emphasizing the importance of pre-defining a “best practice” note format, rather than relying on LLMs to determine this for clinical practice.  [^1] Our code and data are available at https://github.com/hanyin88/llama-clinic.
@@
Hanyin Wang1,2, Chufan Gao2, Bolun Liu1, Qiping Xu1, Guleid Hussein1, Mohamad El Labban1, Kingsley Iheasirim1, Hariprasad Korsapati1, Chuck Outcalt 3, Jimeng Sun2,4 1 Mayo Clinic Health System, 2 School of Computing and Data Science, UIUC 3 Mayo Clinic Rochester, 4 Carle Illinois College of Medicine, UIUC wang.hanyin@mayo.edu, jimeng@illinois.edu
@@ 1 Introduction
Recent advancements in LLMs have transformed the field of natural language processing (NLP). However, the application of LLMs in the medical domain is still in its early stages  He et al. (2023); Zhou et al. (2023b). Proprietary LLMs, such as GPT-4 and Med-PaLM, have demonstrated impressive capabilities in medical knowledge and clinical NLP tasks Nori et al. (2023); Singhal et al. (2023); Van Veen et al. (2023). However, most proprietary LLMs have limited flexibility for domain-specific fine-tuning, primarily due to restricted access to model weights. Additionally, proprietary LLMs raise several concerns pertinent to the healthcare sector, including HIPAA compliance, data security, cost, and transparency of training data Marks and Haupt (2023); Clusmann et al. (2023); Adimi (2024); Michael M. Grynbaum (2023).
@@
The emergence of powerful open-source LLMs has opened up opportunities for domain-specific fine-tuning within the clinical field, yielding promising results Han et al. (2023); Wang et al. (2023); Yunxiang et al. (2023); Wu et al. (2023a); Toma et al. (2023); Chen et al. (2023). However, most research on open-source models has concentrated on medical knowledge injection rather than practical applications in real-world clinical workflow such as clinical note generation.
@@
In this work, we address a practical question clinicians face in their everyday routine: How can we best adapt an open-source LLM for the specific use case of clinical note generation? Clinical note documentation represents a significant burden for healthcare practitioners Ammenwerth and Spötl (2009) and appears to be a natural application for LLMs, given their remarkable generative capabilities Lee et al. (2023). Recent research on LLMs for clinical text summarization found that LLM-generated outputs are preferred over human summaries for their completeness and accuracy Van Veen et al. (2023).
@@
Figure 1: Overview of Study Design. We conducted a comprehensive domain- and task-specific adaptation process for the LLaMA-2-13B model. This process included continued pretraining, supervised fine-tuning, and reinforcement learning from AI and human feedback. Finally, we evaluated our model’s outputs against those created by physicians and Gemini Pro through a blinded expert evaluation. We used Gemini 1.0 Pro as the teacher model in this study.
@@
The 2023 ACL ClinicalNLP and CLEFImage workshops explored the generation of clinical notes from patient-doctor conversations using the newly released Ambient Clinical Intelligence Benchmark (ACI-BENCH) dataset Abacha et al. (2023); Yim et al. (2023a). The most notable results were achieved using GPT-4 along with few-shot in-context learning Giorgi et al. (2023); Van Veen et al. (2023). However, these initial exploratory studies for clinical note generation with LLMs revealed significant limitations:
@@
  • Issue with metrics: The evaluation relied solely on automatic metrics of lexical similarity to the “reference” notes, such as ROUGE scores, which may not accurately reflect human preferences Moramarco et al. (2022); Van Veen et al. (2023).
@@
  • Limited fine-tuning datasets: Previous studies conducted only limited supervised fine-tuning (SFT) of open-source LLMs using small training datasets, thereby not fully exploring their potential for domain- and task-specific adaptation Abacha et al. (2023); Yim et al. (2023a); Van Veen et al. (2023).
@@
    Clinical LLM fine-tuning playbook: Our relatively compact, open-source LLM demonstrated expert-level performance, achieving the same quality as physician-authored notes in the “Assessment and Plan" section. We provide a comprehensive guide for healthcare organizations to fine-tune local LLMs on their own data.
@@
    Key considerations for future clinical note-generation tasks: We underscored the importance of pre-defining a “best practice” note format, rather than relying on LLMs to determine this for clinical practice. We also recommend a workflow for physicians to work with AI-generated notes in real-world practice, ensuring physicians’ full oversight and ultimate accountability for the AI-generated content.
@@ 2.1 Problem Formation
Given a recorded dialogue from a patient-doctor clinic encounter, we task LLMs to generate a high-quality outpatient note akin to one written by a clinician. This scenario is becoming increasingly prevalent due to the rising popularity of ambient AI scribe products Yim et al. (2023b); Barr et al. (2024). Our research focus is on generating the “Subjective” and “Assessment and Plan” sections of outpatient notes. This decision is based on feedback from our physician coauthors, which indicates that discussing all details of the “Objective” section, such as physical examination results, is impractical during real-world clinical encounters. Furthermore, much of the objective data, including lab and imaging results, are directly integrated into Electronic Medical Records (EMR), making clinical notes generation for those sections easy and sometimes unnecessary. Our research is also related to prior work on generating clinical assessments within notes Yang and Yu (2020) and on generating doctor-patient conversations Wang et al. (2024c).
@@ 2.2 ACI-BENCH
ACI-BENCH represents the largest clinic dialogue-note dataset publicly available to date, comprising 207 cases Yim et al. (2023b). The dataset’s dialogues were crafted by a team with medical expertise, and its clinical notes were initially generated using an automatic note generation system, then reviewed and revised by domain experts Yim et al. (2023b). The ACI-BENCH dataset was previously utilized to benchmark the performance of outpatient note generation systems using automatic metrics that evaluate lexical similarity Abacha et al. (2023); Yim et al. (2023a). Importantly, we observed notable variation in the format, style, and quality of the “reference notes” within ACI-BENCH, especially in the section of “Assessment and Plan”. While this diversity mirrors the reality of clinical practice, where different doctors may produce vastly different notes, it presents a challenge to use these notes as a “gold standard” for training an LLM to replicate.
@@ 2.3 Distilled DPO
Given the significant time and financial costs associated with collecting preference data, utilizing pre-collected preference datasets, such as those employed in distilled DPO, is a common practice. Distilled DPO involves generating a collection of responses for each prompt from various LLMs (Figure 2A) Tunstall et al. (2023). These responses are then evaluated by a teacher model (e.g., GPT-4) to provide preference feedback. Applying reinforcement learning with AI feedback (RLAIF) with distilled DPO has yielded encouraging outcomes Tunstall et al. (2023). However, distilled DPO’s reliance on a pre-collected preference dataset renders it suboptimal due to the off-policy and offline characteristics.
@@ 3 DistillDirect
3.1 Comparison of Online vs. Offline and On-Policy vs. Off-Policy Training
@@ 3.2 RLAIF with DistillDirect
In our study, we employed RLAIF with DistillDirect, followed by Reinforcement Learning from Human Feedback (RLHF), which provides several advantages. First, for each iteration of DistillDirect, we used the target policy’s outputs as “reject samples” to inform feedback, ensuring on-policy learning as opposed to the off-policy training in previous work Tunstall et al. (2023). Second, during the RLHF phase, we gathered human preferences on responses generated by the target policy, promoting online and on-policy training (Figure 2C).
@@ 4.1 Experimental Design
Model Selection: We selected Meta’s LLaMA-2-13B and conducted experiments using both the base and chat models Touvron et al. (2023). Gemini 1.0 Pro (hereafter referred to as Gemini Pro) by Google was selected as the teacher model for generating reference notes Team et al. (2023).
Experiments Overview: We demonstrated the experiments pipeline in Figure 1. First, we undertook domain-specific adaptation of LLaMA-2-13B through continued pretraining on discharge summaries from MIMIC-IV. This was followed by task-specific fine-tuning with SFT through instruction tuning. Next, we conducted RLAIF using DistillDirect, our enhanced approach to performing DPO on a distilled dataset. We then selected the model that performed best after SFT and RLAIF, as measured by ROUGE scores against reference notes, for RLHF via DPO. Finally, a panel of physicians conducted a blinded evaluation of the notes authored by our LLaMA-Clinic model, Gemini Pro, and other physicians. All training was performed using low-rank adaptation (LoRA).
@@ 4.2 Dataset and Preprocessing
Modified ACI-BENCH: Given above mentioned limitation of ACI-BENCH, we established a simple yet specific note format, recognized as a “best practice” by a panel of licensed internal medicine physicians, to standardize our training approach. For model training, we retained only the dialogue section from ACI-BENCH and employed Gemini Pro to generate notes based on the “best practice” format, serving as our reference notes. We demonstrated two examples of clinical notes before and after the change in Appendix Figure 7.
Dialogue-G: We created a synthetic dataset of clinical dialogue-note pairs using Gemini Pro. This dataset, named Dialogue-G, comprises 1,291 cases. We first compiled transcribed outpatient notes from the publicly available synthetic MTSamples dataset MTSamples ; Hu et al. (2024) and utilized Gemini Pro to transform these notes into dialogues. Subsequently, we used these dialogues as inputs for Gemini Pro once again, this time to generate clinical notes based on the “best practice” format.
MIMIC-IV: MIMIC-IV encompasses 431,231 unique hospital admissions from Beth Israel Deaconess Medical Center in Boston, Massachusetts Johnson et al. (2023). We utilized discharge summaries from MIMIC-IV for continued pretraining. Notably, the “brief hospital course” section of the discharge summaries is structurally akin to the “assessment and plan” section in outpatient notes. We compiled a subset of discharge summaries with only the “brief hospital course” using methods detailed in Wang et al. (2024a), referred to as Discharge-short. We denoted the complete discharge summaries dataset as Discharge-long.
@@ 4.3 Experiment setup
Dataset Split for SFT, RLAIF and RLHF: We combined the training subsets from ACI-BENCH (dialogue n = 67) and Dialogue-G (dialogue n = 1291), then split this data equally for SFT and RLAIF, stratified by the data source. For RLHF, we utilized dialogues from the training, test2, and test3 subsets of ACI-BENCH (dialogue n = 147).
@@
Physician Reader Study: The three internal medicine physicians engaged in preference data collection were tasked with writing clinical notes based on conversations from the ACI-BENCH test1 subset (dialogue n = 40), adhering to the pre-defined “best practice” format. The physician-authored notes, alongside those generated by Gemini Pro and LLaMA-Clinic, were reviewed by another four physicians, who were not involved in the preference labeling.
@@ 5.1 Analysis of Continued Pretraining
We presented the training loss curve in Figure 3. Across all experiments, the training loss rapidly decreased after the initial few hundred steps, then leveled off, showing minimal improvement thereafter. The lowest training loss achieved with the Discharge-long dataset is approximately 0.9, whereas, with the Discharge-short dataset, it remained around 1.4. The trajectories of training loss were similar for both the chat and base models. When experimenting with various training strategies and hyperparameters, we frequently observed loss spikes that were slow to recover, as shown in Appendix Figure 1. We proceeded with checkpoints that did not exhibit loss spikes for SFT and RLAIF.
@@ 5.2 Analysis of SFT and RLAIF
Model ROUGE-1 ROUGE-2 ROUGE-L ROUGE-LSUM ROUGE-1 ROUGE-2 ROUGE-L ROUGE-LSUM
Baseline: Out-of-Box LLaMA-2 and Medical LLMs
13B 0.0329 0.0033 0.0211 0.0268 0.0100 0.0006 0.0058 0.0085
13B-chat 0.3585 0.1281 0.2103 0.2890 0.4543 0.1778 0.2898 0.4261
Meditron-7B 0.1249 0.0052 0.0678 0.0957 0.0895 0.0032 0.0527 0.0810
LLaMA3-Med42-8B 0.2496 0.0662 0.1392 0.2056 0.2372 0.0577 0.1426 0.2187
MeLLaMA-13B-chat 0.2756 0.1268 0.1925 0.2329 0.2655 0.0964 0.1865 0.2522
SFT + RLAIF Models (No Pretraining)
13B
SFT 0.2813 0.1166 0.1975 0.2248 0.2977 0.1059 0.1963 0.2677
SFT + DistillDirect 0.4994 0.2633 0.3425 0.3964 0.4941 0.2398 0.3476 0.4650
13B-chat
SFT 0.2874 0.1179 0.2022 0.2285 0.3301 0.1191 0.2231 0.2992
SFT + DistillDirect 0.4783 0.2472 0.3252 0.3738 0.4893 0.2411 0.3674 0.4599
Continued Pretraining + SFT + RLAIF Models
13B-long
CP 0.1612 0.0139 0.0754 0.1378 0.1289 0.0108 0.0598 0.1178
CP + SFT 0.2525 0.1034 0.1714 0.2044 0.2565 0.0775 0.1651 0.2276
CP + SFT + DistillDirect 0.4494 0.2471 0.3224 0.3618 0.4578 0.2086 0.3272 0.4241
13B-short
CP 0.1488 0.0106 0.0712 0.1149 0.0899 0.0060 0.0449 0.0806
CP + SFT 0.2463 0.0858 0.1638 0.1951 0.2277 0.0619 0.1396 0.1960
CP + SFT + DistillDirect 0.4775 0.2561 0.3452 0.3842 0.4956 0.2328 0.3526 0.4663
13B-chat-long
CP 0.1708 0.0151 0.0792 0.1321 0.1536 0.0130 0.0746 0.1424
CP + SFT 0.3463 0.1591 0.2525 0.2837 0.3620 0.1237 0.2401 0.3290
CP + SFT + DistillDirect 0.4601 0.2504 0.3325 0.3773 0.4662 0.2395 0.3484 0.4498
13B-chat-short
CP 0.1520 0.0116 0.0725 0.1085 0.1162 0.0107 0.0630 0.1032
CP + SFT 0.3475 0.1310 0.2230 0.2711 0.3055 0.1036 0.2009 0.2824
CP + SFT + DistillDirect 0.4878 0.2613 0.3410 0.3883 0.5182 0.2689 0.3933 0.4915
@@
Table 1: ROUGE Scores Following Continued Pretraining (CP), Supervised Fine-Tuning (SFT), and Reinforcement Learning from AI Feedback (RLAIF). All evaluations were conducted on the validation subset of the modified ACI-BENCH. The “13B” models represent the LLaMA-2 base models, while the “13B-chat” models correspond to the LLaMA-2 chat models. Models with the suffix “long” were pretrained using the Discharge-long dataset, and those with the suffix “short” were pretrained using the Discharge-short dataset. Bolded scores represent the top two scores for each metric. Continued pretraining with MIMIC-IV discharge summaries initially resulted in performance deterioration, whereas SFT and DistillDirect significantly improved the performance across all models.
@@
Since the primary objective of SFT and RLAIF is to align LLaMA-2’s output with Gemini Pro, we evaluated ROUGE scores—a measure of lexical similarity—against reference notes generated by Gemini Pro. We reported ROUGE scores Lin (2004) post-SFT and RLAIF, alongside those from baseline models in Table 1. Interestingly, at baseline, medical LLMs fine-tuned on biomedical literature or clinical notes Chen et al. (2023); Christophe et al. (2024); Xie et al. (2024) underperform compared to the vanilla LLaMA-2 chat model. This suggests that generating notes aligned with our “best practice” format presents an out-of-distribution challenge for medical LLMs. As expected, continued pretraining with MIMIC-IV discharge summaries compromised the chat model’s capacity to follow instructions. SFT notably enhanced performance, particularly for the chat model compared to the base model. The application of RLAIF with DistillDirect significantly boosted performance across all models. In many cases, the gains achieved through RLAIF surpassed those from SFT, as reflected by the delta improvement in ROUGE scores (e.g., 13B-short). Notably, our training with DistillDirect frequently encountered instability—a well-known challenge in reinforcement learning Ding and Dong (2020). With carefully selected training hyperparameters, DistillDirect could achieve performance improvements without overfitting, despite the limited training data and repetitive use of the same prompts in each training cycle (see Appendix Table 4). We have detailed our experiments, including ablation studies, to find a stable training setup in Appendix Method A.3.
@@
When comparing models pretrained with Discharge-long and Discharge-short, the latter consistently achieved higher ROUGE scores, despite exhibiting higher training loss during the pretraining phase. The chat model pretrained with the Discharge-short dataset (13B-chat-short) emerged as the top performer for the final RLHF stage. Interestingly, the model trained with SFT and RLAIF, without continued pretraining, demonstrated strong performance, including the highest score for “Subjective.” This raises questions about the benefits of continued pretraining, particularly considering its substantial computational demands (Appendix Figure 6A).